OpenStudy (anonymous):

f(x)=25x^2+40x+16, find the x- and y intercepts.

OpenStudy (anonymous):

(5x+4)(5x+4)

OpenStudy (anonymous):

don't forget the y-intercept at (0,16)

OpenStudy (anonymous):

(5x+4)^2 x=-4/5 plug x=0 find if x=0 than y intercept is 16

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

plug y=0 than find x

OpenStudy (anonymous):

T Y